OpenStudy (anonymous):

solve for g q= 3g+3z g= ?

OpenStudy (anonymous):

g = (q - 3z)/3 or g = q/3 - z

OpenStudy (anonymous):

or?

OpenStudy (anonymous):

what would g=

OpenStudy (anonymous):

g is a dependent variable of q,z. You can say that g is a function of q,z.

OpenStudy (anonymous):

okay